# Gamma: Toward Generic Image Assessment with Mixture of Assessment Experts

### v1

@@ Abstract
Image assessment aims to evaluate the quality and aesthetics of images and has been applied across various scenarios, such as natural and AIGC scenes. Existing methods mostly address these sub-tasks or scenes individually. While some works attempt to develop unified image assessment models, they have struggled to achieve satisfactory performance or cover a broad spectrum of assessment scenarios. In this paper, we present Gamma, a Generic imAge assessMent model using Mixture of Assessment Experts, which can effectively assess images from diverse scenes through mixed-dataset training. Achieving unified training in image assessment presents significant challenges due to annotation biases across different datasets. To address this issue, we first propose a Mixture of Assessment Experts (MoAE) module, which employs shared and adaptive experts to dynamically learn common and specific knowledge for different datasets, respectively. In addition, we introduce a Scene-based Differential Prompt (SDP) strategy, which uses scene-specific prompts to provide prior knowledge and guidance during the learning process, further boosting adaptation for various scenes. Our Gamma model is trained and evaluated on 12 datasets spanning 6 image assessment scenarios. Extensive experiments show that our unified Gamma outperforms other state-of-the-art mixed-training methods by significant margins while covering more scenes.
Machine Learning, ICML
1 Introduction
@@
Image assessment is a long-standing research topic in the field of image processing, primarily comprising two tasks: Image Quality Assessment (IQA) and Image Aesthetic Assessment (IAA). These tasks require models to automatically evaluate the visual quality and aesthetic appeal of images, respectively. They have broad applications in various real-world scenarios, such as guiding image dehazing (Zhao et al., 2021), selecting high-quality images in a data engine (Rombach et al., 2022), serving as tools in an agentic system (Yang et al., 2024), or acting as reward models when aligning image generative models with human feedback (Liang et al., 2024).
@@
Due to differences in image content and application scenarios, image assessment has spawned many sub-tasks, such as Natural-IQA for natural images, Face-IQA for facial images, AIGC-IQA for generated images, and IAA. Accordingly, numerous methods (Ke et al., 2021; He et al., 2022; Su et al., 2023b) have been proposed to address these specific tasks. However, these models often struggle to apply directly to other scenes or typically require task-specific fine-tuning on a given dataset. As illustrated in Figure 1, it is challenging for DEIQT (Qin et al., 2023) to transfer to other datasets without fine-tuning. This limitation prevents image assessment models from being widely applicable, e.g., IQA models are needed to assess facial, artistic, and natural images in the AIGC scene. Hence, there is an urgent need for a model that can effectively handle a variety of scenarios.
@@
Figure 1: “DEIQT Finetune” is trained and tested on the same dataset. “DEIQT ZS” directly assesses images using the model trained on other datasets, which perform poorly. PromptIQA and our Gamma are trained on mixed datasets. Our Gamma performs well on multiple datasets simultaneously and even surpass the task-specific method DETQT.
Figure 2: Images with similar MOS labels from different datasets exhibit drastically different perceptual quality. It is not hard to observe that image (a) has clearly superior quality than the other three. Zoom in for a better view.
@@
To mitigate this issue, some approaches attempt to combine many datasets from various assessment tasks to train a general image assessment model. For instance, UNIQUE (Zhang et al., 2021) and LIQE (Zhang et al., 2023) utilize multiple authentic and synthetic natural IQA datasets for mixed training, but they focus only on Natural-IQA. Q-Align (Wu et al., 2023) uses a large language model with billions of parameters to unify IQA and IAA tasks, but it has a low inference speed and focuses solely on natural images. Additionally, PromptIQA (Chen et al., 2024b) employs image-score pairs as prompts for quality predictions, but it is inflexible as it requires multiple images as references during inference. These methods, however, fail to achieve competitive performance compared to models trained on specific datasets and cover a broad range of scenes. Thus, it is imperative to develop a foundational image assessment model capable of evaluating images from various scenes.
@@
To this end, we present Gamma, a Generic imAge assessMent model using Mixture of Assessment experts, to achieve unified image assessment across multiple datasets effectively. We found that the primary challenge in mixed-dataset training is the mean opinion score (MOS) bias between different datasets, e.g., images with similar MOS may exhibit different perceptual qualities across various datasets. As shown in Figure 2, samples from KonIQ (Hosu et al., 2020), UWIQA (Yang et al., 2021a), and SPAQ (Fang et al., 2020) show obvious differences in perceptual quality despite being labeled with similar MOS. To address this challenge, we introduce a novel Mixture of Assessment Experts (MoAE) module in our Gamma model. MoAE consists of two types of experts: shared experts and adaptive experts. The shared experts are employed throughout to learn dataset-shared knowledge, while the adaptive experts are dynamically activated to varying degrees to learn dataset-specific knowledge. Additionally, an image-based router modulates the contributions of each adaptive expert. This strategy allows the model to capture common features while flexibly adjusting representative features for different datasets. Compared with general Mixture of Experts (MoE) (Shazeer et al., 2017) and Lora-based MoE (Liu et al., 2024), we equip the MoAE module only in the rear blocks instead of all blocks, making it more efficient. Secondly, we propose a Scene-based Differential Prompt (SDP), which uses different prompts for different datasets according to their scenes. This strategy provides scene-specific knowledge for representation learning of different datasets, guiding the mixed-dataset training process.
@@
2 Related Work
2.1 Image Assessment
Image Assessment mainly includes two subtasks: Image Quality Assessment (IQA) and Image Aesthetic Assessment (IAA). In the deep learning era, these two tasks have achieved significant breakthroughs. For the IQA task, researchers develop various advanced techniques to improve performance, including multi-level feature aggregation (Li et al., 2018; Chen et al., 2024a; Xu et al., 2024; Zhang et al., 2018; Mittal et al., 2012b; Ying et al., 2020), adaptive convolution (Su et al., 2020), transformer methods (Ke et al., 2021; Qin et al., 2023), vision-language models (VLMs) (Wang et al., 2023; Zhang et al., 2023) and large language models (LLM) (You et al., 2023). Moreover, besides the natural image assessment, these are various IQA methods for other scenes, such as face IQA (Ou et al., 2021; Su et al., 2023b; Jo et al., 2023), AIGC IQA (Yuan et al., 2023), underwater IQA (Yang et al., 2021b; Guo et al., 2023; Yang & Sowmya, 2015; Liu et al., 2023). For the IAA task, numerous methods have also been proposed to improve performance, including loss function (Talebi & Milanfar, 2018), novel transformer architecture (Tu et al., 2022), multi-level features (Hosu et al., 2019), theme information (He et al., 2022; Li et al., 2023b) and multimodal pre-training (Ke et al., 2023).
2.2 Mixed Training for Image Assessment
@@ 2.3 Mixture of Experts
Figure 3: (a) The architecture of Gamma: It consists of a visual encoder $\mathcal{V}$ and text encoder $\mathcal{T}$; We add the Mixture of Assessment Experts (MoAE) to the last few layers of both encoders. We introduce a Scene-based Differential Prompt (SDP) to prompt images from different scenes (See Section 3.4 for details). (b) The MoAE module: It involves one shared expert ($E^{\mathrm{shared}}$) and several adaptive experts (from $E_{1}^{\mathrm{adaptive}}$ to $E_{n}^{\mathrm{adaptive}}$). We employ a router to adaptively activate the adaptive experts. We then use a learnable factor $\sigma$ to merge the features of two type of experts.
3 Method
3.1 Preliminary
As a foundational vision-language model (VLM), CLIP (Radford et al., 2021) has shown significant promise in supporting a wide array of vision tasks. Specifically, CLIP is composed of a transformer-based visual encoder $\mathcal{V}$ and a text encoder $\mathcal{T}$, which generate aligned visual representations $I$ and text representations $T$ for each image-text pair. Utilizing these features, we can compute cosine similarity scores between image and text pairs across different domains or tasks to perform task-specific predictions, including image assessment tasks. Recently, to enhance CLIP’s capabilities in the field of image assessment, UniQA (Zhou et al., 2024b) fine-tuned CLIP on large-scale synthetic and authentic image-text datasets focused on image quality and aesthetics. This approach demonstrates excellent performance on both IQA and IAA tasks after task-specific fine-tuning. However, the model lacks foundational applicability across various image assessment scenarios without fine-tuning. Building on the unified training pipeline proposed in UniQA, we propose two approaches to confront MOS bias present in different datasets and develop a foundational image assessment model. In the following, we will provide a detailed exposition of its components.
3.2 Overview of Gamma
@@
$\mathrm{q}=\sum_{k=1}^{5}\mathrm{C}_{k}\mathrm{Softmax}({I^{\prime}}^{\top}T_{k}/\tau),I^{\prime}=\mathrm{Adapter}(I),$ (1)
@@
3.3 Mixture of Assessment Experts
To develop a unified and generic image assessment model, we aim to combine multiple image assessment datasets for joint training. Unfortunately, the mean opinion score (MOS) introduces significant biases across different datasets, which hinders joint training. To address this challenge, we propose the MoAE module, where several experts are employed to learn the diverse biases of different datasets. As shown in Figure 3 (b), the proposed MoAE module includes a shared assessment expert ($E^{\mathrm{shared}}$) to learn common knowledge of image assessment and several adaptive assessment experts (from $E_{1}^{\mathrm{adaptive}}$ to $E_{n}^{\mathrm{adaptive}}$) to dynamically learn dataset-specific knowledge.
The Shared Assessment Expert. The shared assessment expert $E^{\mathrm{shared}}$ inherits the image assessment capabilities of the original CLIP model by reusing its weights. This expert remains frozen during training to ensure that the learned world knowledge is retained. Thus, the model can capture common representations across various contexts and maintain its original multi-modal capabilities. Given an input hidden state $x\in\mathbb{R}^{d}$, the output of the shared assessment expert is:
$y^{\mathrm{shared}}=E^{\mathrm{shared}}(x),$ (2)
@@ 3.3 Mixture of Assessment Experts
Figure 4: Visual examples from different datasets, which include natural images, underwater images, face images, AI-generated images, and etc. The first value is the prediction score and the second value is the ground-truth MOS. Our Gamma can accurately evaluate images from different scenes, demonstrating the generalization and effectiveness. All images are resized for better visibility.
@@
$y^{\mathrm{adaptive}}$ $=\sum_{i=1}^{n}G(x)_{i}E_{i}^{\mathrm{adaptive}}(x),$ (3)
@@
Here, the router $G(\cdot)$ is a linear transformation for the input feature $x$; $W\in\mathbb{R}^{n\times d}$ is the transformation matrix. To avoid unreasonable weights, we utilize a $\mathrm{Softmax}$ operator to normalize the contribution weights. This ensures that the model can learn dataset-specific knowledge efficiently. Notably, we initialize these experts with the weights of pre-trained CLIP to equip them with image assessment capabilities.
@@
$y^{\mathrm{MoAE}}=y^{\mathrm{shared}}+\sigma\cdot y^{\mathrm{adaptive}}.\$ (4)
The $\sigma$ factor is zero-initialized so that the visual and text encoders can generate aligned features at the beginning. In practice, we freeze the shared experts and set the adaptive experts to be tunable only. This approach maintains parameter efficiency during mixed training and preserves the multi-modal capabilities of the original model.
We incorporate the MoAE module into the last $K$ layers of both visual and text encoders, as shown in Figure 3. This strategy makes our method both effective and efficient. The visual feature extraction process can be formulated as:
$I_{i}$ $=\bm{\mathcal{V}}_{i}(I_{i-1}),\quad i=1,2,\ldots,L-K$ (5)
@@
3.4 Scene-based differential prompt
To facilitate scene-guided learning, we introduce a Scene-based Differential Prompt (SDP) to help the model acquire scene-specific knowledge from different datasets. We utilize 12 datasets spanning 6 image assessment scenarios, including synthetic distortion nature IQA, authentic distortion nature IQA, face IQA, AIGC IQA, underwater IQA, and IAA, for mixed training. We categorize these datasets into five groups based on their scenes: natural quality, AI-generated quality, underwater quality, face quality, and natural aesthetics. Specifically, for the face quality assessment dataset, we use prompts such as face bad-quality, face poor-quality, face fair-quality, face good-quality, face perfect-quality appended to the word image. For more details on these prompts, please refer to Appendix A.4.
4 Experiments
4.1 Datasets
We utilize 12 datasets for unified training and testing, encompassing both image quality and aesthetic assessment tasks. For the IQA task, five different assessment scenarios are included: synthetic distortion nature IQA (SDN-IQA), authentic distortion nature IQA (ADN-IQA), face IQA (F-IQA), AIGC IQA (AG-IQA), and underwater IQA (U-IQA). For the IAA task, we use two classical benchmarks, AVA and AADB. In addition, we use two rare datasets to verify the generalization ability of the model, i.e., exBeDDE and ECIQAD. The exBeDDE is a dehazed image quality assessment (D-IQA) dataset, and the ECIQAD is an enhanced colonoscopy image quality assessment (EC-IQA) dataset. Details about these datasets is provided in Table 9.
Table 1: Comparison with SOTA task-specific and mixed-training models on 12 datasets for 6 image assessment tasks. “Gamma” and “Gamma-T” denote the mixed-training and task-specific models, respectively. Gamma uses the Scene-based Differential Prompt (SDP) for training and testing. indicates that we retrain the model with the same data split as ours.
Task Synthetic Distortion Nature IQA (SDN-IQA) Authentic Distortion Nature IQA (ADN-IQA)
@@ 4.1 Datasets
SDD-FIQA 0.602 0.649 DBCNN 0.821 0.876 FDUM 0.694 0.689 MaxViT 0.708 0.745 MUSIQ 0.706 0.712
Task Specific IFQA 0.697 0.722 HyperNet 0.836 0.890 UCIQE 0.627 0.626 TANet 0.758 0.765 TANet 0.738 0.737
@@
GPFIQA 0.964 0.965 PSCR 0.850 0.906 UIQI 0.742 0.741 UniQA 0.776 0.776 UniQA 0.786 0.787
@@
Table 2: The effect of Mixture of Assessment Experts (MoAE) and Scene-based Differential Prompt (SDP).

### v2

@@ Abstract.
Image assessment aims to evaluate the quality and aesthetics of images and has been applied across various scenarios, such as natural and AIGC scenes. Existing methods mostly address these sub-tasks or scenes individually. While some works attempt to develop unified image assessment models, they have struggled to achieve satisfactory performance or cover a broad spectrum of assessment scenarios. In this paper, we present Gamma, a Generic imAge assessMent model using Mixture of Assessment Experts, which can effectively assess images from diverse scenes through mixed-dataset training. Achieving unified training in image assessment presents significant challenges due to annotation biases across different datasets. To address this issue, we first propose a Mixture of Assessment Experts (MoAE) module, which employs shared and adaptive experts to dynamically learn common and specific knowledge for different datasets, respectively. In addition, we introduce a Scene-based Differential Prompt (SDP) strategy, which uses scene-specific prompts to provide prior knowledge and guidance during the learning process, further boosting adaptation for various scenes. Our Gamma model is trained and evaluated on 12 datasets spanning 6 image assessment scenarios. Extensive experiments show that our unified Gamma outperforms other state-of-the-art mixed-training methods by significant margins while covering more scenes. Codes are available at https://github.com/zht8506/Gamma.
Generic Image Assessment, Unified Training, Mixture of Experts
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33nd ACM International Conference on Multimedia; October 27-31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33nd ACM International Conference on Multimedia (MM ’25), October 27-31, 2025, Dublin, Irelanddoi: XXXX.XXXXisbn: ccs: Computing methodologies Computer vision
1. Introduction
@@
Image assessment is a long-standing research topic in the field of image processing, primarily comprising two tasks: Image Quality Assessment (IQA) and Image Aesthetic Assessment (IAA). These tasks require models to automatically evaluate the visual quality and aesthetic appeal of images, respectively. They have broad applications in various real-world scenarios, such as guiding image dehazing (Zhao et al., 2021), selecting high-quality images in a data engine (Rombach et al., 2022), serving as tools in an agentic system (Yang et al., 2024a), or acting as reward models when aligning image generative models with human feedback (Liang et al., 2024).
@@
Due to differences in image content and application scenarios, image assessment has spawned many sub-tasks, such as Natural-IQA for natural images, Face-IQA for facial images, AIGC-IQA for generated images, and IAA. Accordingly, numerous methods (Ke et al., 2021; He et al., 2022; Su et al., 2023a) have been proposed to address these specific tasks. However, these models often struggle to apply directly to other scenes or typically require task-specific fine-tuning on a given dataset. As illustrated in Figure 1, it is challenging for DEIQT (Qin et al., 2023) to transfer to other datasets without fine-tuning. This limitation prevents image assessment models from being widely applicable, e.g., IQA models are needed to assess facial, artistic, and natural images in the AIGC scene (Rombach et al., 2022; Yang et al., 2024a; Liang et al., 2024). Hence, there is an urgent need for a model that can effectively handle a variety of scenarios.
@@
Figure 1. “DEIQT Finetune” is trained and tested on the same dataset. “DEIQT ZS” directly assesses images using the model trained on other datasets, which perform poorly. PromptIQA and our Gamma are trained on mixed datasets. Our Gamma performs well on multiple datasets simultaneously and even surpass the task-specific method DETQT.
@@
To mitigate this issue, some approaches attempt to combine many datasets from various assessment tasks to train a general image assessment model. For instance, UNIQUE (Zhang et al., 2021) and LIQE (Zhang et al., 2023) utilize multiple authentic and synthetic natural IQA datasets for mixed training, but they focus only on natural image. Q-Align (Wu et al., 2023) uses a large language model with billions of parameters to unify IQA and IAA tasks, but it has a low inference speed and still solely focuses on natural images. PromptIQA (Chen et al., 2024b) employs image-score pairs as prompts for quality predictions, but it is inflexible as it requires multiple images as references during inference. These methods fail to cover a broad range of scenes and achieve competitive performance compared to models trained on specific datasets. Thus, it is imperative to develop an effective and generic image assessment model capable of evaluating images from various scenes.
@@
To this end, we present Gamma, a Generic imAge assessMent model using Mixture of Assessment experts, to achieve unified image assessment across multiple scenes effectively. We found that the primary challenge in mixed-dataset training is the mean opinion score (MOS) bias between different datasets, e.g., images with similar MOS may exhibit different perceptual qualities across various datasets. As shown in Figure 2, samples from KonIQ (Hosu et al., 2020), UWIQA (Yang et al., 2021a), and SPAQ (Fang et al., 2020) show obvious differences in perceptual quality despite being labeled with similar MOS. To address this challenge, we introduce a novel Mixture of Assessment Experts (MoAE) module in our Gamma model. It consists of two types of experts: shared experts and adaptive experts. The shared experts are employed throughout to learn dataset-shared knowledge, while the adaptive experts are dynamically activated to varying degrees to learn dataset-specific knowledge. Additionally, an image-based router modulates the contributions of each adaptive expert. The MoAE allows the model to capture common features while flexibly adjusting representative features for different datasets. Compared with general Mixture of Experts (MoE) (Shazeer et al., 2017) and LoRA-based MoE (Liu et al., 2024b), we equip the MoAE only in the rear blocks instead of all blocks, making it more efficient. Secondly, we propose a Scene-based Differential Prompt (SDP), which uses different prompts for different datasets according to their scenes. This strategy provides scene-specific knowledge for representation learning of different datasets, guiding the mixed-dataset training process.
Figure 2. Images with similar MOS labels from different datasets exhibit drastically different perceptual quality. It is not hard to observe that image (a) has clearly superior quality than the other three. Zoom in for a better view.
@@
2. Related Work
2.1. Image Assessment
Image Assessment mainly includes two subtasks: Image Quality Assessment (IQA) and Image Aesthetic Assessment (IAA). In the deep learning era, these two tasks have achieved significant breakthroughs. For the IQA task, researchers develop various advanced techniques to improve performance, including feature and model structure design (Ying et al., 2020; Zhang et al., 2018), multi-level feature aggregation (Li et al., 2018; Chen et al., 2024a; Xu et al., 2024), adaptive convolution (Su et al., 2020), transformer methods (Ke et al., 2021; Qin et al., 2023), vision language models (VLMs) (Wang et al., 2023; Zhang et al., 2023) and large language models (LLM) (You et al., 2023; Wu et al., 2023). Moreover, besides the natural image assessment, these are various IQA methods for other scenes, such as face IQA (Ou et al., 2021; Su et al., 2023a; Jo et al., 2023), AIGC IQA (Yuan et al., 2023), underwater IQA (Yang et al., 2021b; Guo et al., 2023; Yang and Sowmya, 2015; Liu et al., 2023). For the IAA task, numerous methods have also been proposed to improve performance, including loss function (Talebi and Milanfar, 2018), novel transformer architecture (Tu et al., 2022), multi-level features (Hosu et al., 2019), theme information (He et al., 2022; Li et al., 2023a) and multimodal pre-training (Ke et al., 2023).
Figure 3. (a) The architecture of Gamma: It consists of a visual encoder $\mathcal{V}$ and text encoder $\mathcal{T}$; We add the Mixture of Assessment Experts (MoAE) to the last few layers of both encoders. We introduce a Scene-based Differential Prompt (SDP) to prompt images from different scenes (See Section 3.4 for details). (b) The MoAE module: It involves one shared expert ($E^{\mathrm{shared}}$) and several adaptive experts (from $E_{1}^{\mathrm{adaptive}}$ to $E_{n}^{\mathrm{adaptive}}$). We employ a router to adaptively activate the adaptive experts. We then use a learnable factor $\sigma$ to merge the features of two type of experts.
2.2. Mixed Training for Image Assessment
@@ 2.3. Mixture of Experts
3. Method
3.1. Preliminary
As a foundational vision-language model (VLM), CLIP (Radford et al., 2021) has shown significant promise in supporting a wide array of vision tasks, including IQA and IAA tasks (Wang et al., 2023; Zhang et al., 2024; Pan et al., 2024). Specifically, CLIP is composed of a visual encoder $\mathcal{V}$ and a text encoder $\mathcal{T}$, which generate aligned visual representations $I$ and text representations $T$ for each image-text pair. Utilizing these features, we can compute cosine similarity scores between image and text pairs across different domains or tasks to perform task-specific predictions, including image assessment tasks. Recently, to enhance CLIP’s capabilities in the field of image assessment, UniQA (Zhou et al., 2024b) fine-tuned CLIP on large-scale synthetic and authentic image-text datasets focused on image quality and aesthetics. This approach demonstrates excellent performance on both IQA and IAA tasks after task-specific fine-tuning. However, the model lacks foundational applicability across various image assessment scenarios without fine-tuning. Building on the unified training pipeline proposed in UniQA, we propose two approaches to confront MOS bias present in different datasets and develop a foundational image assessment model. In the following, we will provide a detailed exposition of its components.
Figure 4. Visual examples from different datasets, which include natural images, underwater images, face images, AIGC images, and etc. The first value is the prediction score and the second value is the ground-truth MOS. Our Gamma accurately evaluates images from different scenes, demonstrating the generalization and effectiveness. All images are resized for better visibility.
3.2. Overview of Gamma
@@
(1) $\mathrm{q}=\sum_{k=1}^{5}\mathrm{C}_{k}\mathrm{Softmax}({I^{\prime}}^{\top}T_{k}/\tau),I^{\prime}=\mathrm{Adapter}(I),$
@@
3.3. Mixture of Assessment Experts
To develop a unified and generic image assessment model, we aim to combine multiple image assessment datasets for joint training. Unfortunately, the mean opinion score (MOS) introduces significant biases across different datasets, which hinders joint training. To address this challenge, we propose the MoAE module, where several experts are employed to learn the diverse biases of different datasets. As shown in Figure 3 (b), the MoAE includes a shared assessment expert ($E^{\mathrm{shared}}$) to learn common knowledge of image assessment and several adaptive assessment experts (from $E_{1}^{\mathrm{adaptive}}$ to $E_{n}^{\mathrm{adaptive}}$) to dynamically learn dataset-specific knowledge.
The Shared Assessment Expert. The shared assessment expert $E^{\mathrm{shared}}$ inherits the image assessment capabilities of the original multimodal model (e.g., CLIP (Radford et al., 2021) or UniQA (Zhou et al., 2024b)) by reusing its weights. This expert remains frozen during training to ensure that the learned world knowledge is retained. Thus, the model can capture common representations across various contexts and maintain its original multi-modal capabilities. Given an input hidden state $x\in\mathbb{R}^{d}$, the output of the shared assessment expert is:
(2) $y^{\mathrm{shared}}=E^{\mathrm{shared}}(x),$
@@ 3.3. Mixture of Assessment Experts
(3) $y^{\mathrm{adaptive}}$ $=\sum_{i=1}^{n}G(x)_{i}E_{i}^{\mathrm{adaptive}}(x),$
@@
Here, the router $G(\cdot)$ is a linear transformation for the input feature $x$; $W\in\mathbb{R}^{n\times d}$ is the transformation matrix. We utilize a $\mathrm{Softmax}$ operator to normalize the contribution weights. Notably, we initialize these experts with the weights of pre-trained multimodal model to equip them with preliminary image assessment capabilities.
@@
(4) $y^{\mathrm{MoAE}}=y^{\mathrm{shared}}+\sigma\cdot y^{\mathrm{adaptive}}.\$
The $\sigma$ factor is zero-initialized so that the visual and text encoders can generate aligned features at the beginning of training. We freeze the shared experts and set the adaptive experts to be tunable only. This approach preserves the learned knowledge of the original model and improves training efficiency.
Inspired by (Yang et al., 2024b), which points out that higher layers are easier to tune for downstream tasks and freezing lower layers preserves more generalizable knowledge, we incorporate the MoAE into the last $K$ layers of both visual and text encoders, as shown in Figure 3. This strategy makes our method both effective and efficient. The visual feature extraction process can be formulated as:
(5) $I_{i}$ $=\bm{\mathcal{V}}_{i}(I_{i-1}),\quad i=1,2,\ldots,L-K$
@@
Table 1. Comparison with SOTA task-specific and mixed-training models on 12 datasets for 6 image assessment tasks. “Gamma” and “Gamma-T” denote the mixed-training and task-specific models, respectively. Gamma uses the Scene-based Differential Prompt (SDP) for training and testing. indicates that we retrain the model with the same data split as ours.
Task Synthetic Distortion Natural IQA (SDN-IQA) Authentic Distortion Natural IQA (ADN-IQA)
@@
IFQA 0.960 0.960 DBCNN 0.821 0.876 FDUM 0.694 0.689 MaxViT 0.708 0.745 MUSIQ 0.706 0.712
Task Specific StyleGAN 0.967 0.968 HyperNet 0.836 0.890 UCIQE 0.627 0.626 TANet 0.758 0.765 TANet 0.738 0.737
@@
CR-FIQA 0.959 0.960 MUSIQ 0.826 0.866 UIQI 0.742 0.741 ELTA 0.764 0.777 ELTA 0.760 0.772
GPFIQA 0.964 0.965 PSCR 0.850 0.906 HPUIQA 0.796 0.790 UniQA 0.776 0.776 UniQA 0.786 0.787
@@
3.4. Scene-based differential prompt
To facilitate scene-guided learning, we introduce a Scene-based Differential Prompt (SDP) to help the model acquire scene-specific knowledge from different datasets. We utilize 12 datasets spanning 6 image assessment scenarios, including synthetic distortion Natural IQA, authentic distortion Natural IQA, face IQA, AIGC IQA, underwater IQA, and IAA, for mixed training. We categorize these datasets into five groups based on their scenes: natural quality, AI-generated quality, underwater quality, face quality, and natural aesthetics. Specifically, for the face quality assessment dataset, we use prompts such as face bad-quality, face poor-quality, face fair-quality, face good-quality, face perfect-quality appended to the word image. Then, the text prompt is sent to the text encoder of the Gamma to obtain text features. Text features will be combined with image features for quality prediction (Section 3.2). For more details on these prompts, please refer to Supplementary Materials.
4. Experiments
4.1. Datasets
We utilize 12 datasets for unified training and testing, encompassing both IQA and IAA tasks. For the IQA task, five different assessment scenarios are included: synthetic distortion natural IQA (SDN-IQA), authentic distortion natural IQA (ADN-IQA), face IQA (F-IQA), AIGC IQA (AG-IQA), and underwater IQA (U-IQA). For the IAA task, we use two classical benchmarks, AVA and AADB. In addition, we use two rare datasets to verify the generalization ability of the model, i.e., exBeDDE and ECIQAD. The exBeDDE is a dehazed image quality assessment (D-IQA) dataset, and the ECIQAD is an enhanced colonoscopy image quality assessment (EC-IQA) dataset. Details about these datasets is provided in supplementary materials.
4.2. Implementation Details
Following the settings in (Su et al., 2020; Ke et al., 2021), we randomly divide each dataset into 80% for training and 20% for testing. The training dataset is a mixture of the training sets of each dataset and we test Gamma on each test data separately. This process is repeated 10 times, and the median of the 10 scores is reported as the final score. We use the pre-trained weight of UniQA (Zhou et al., 2024b), which uses CLIP-B/16 as multimodal encoder. We freeze the CLIP visual and text encoders, training only the MoAE module and adapter. For the unified training, we train the model for 10 epochs with a batch size of 8. The initial learning rate is set to 2e-5. We normalize the MOS/DMOS scale to [0, 1] and utilize Adam optimizer (Kingma, 2014) and MSE loss to optimize the model. For the task-specific training, we use different training settings according to the task and size of datasets. More training details are provided in the supplementary materials.
Table 2. The effect of Mixture of Assessment Experts (MoAE) and Scene-based Differential Prompt (SDP).